# Dark matter admixed neutron stars with a realistic nuclear equation of state from chiral nuclear interactions

### v1

@@ VII Summary
In this paper we have studied the consequences of the presence of DM on the structural properties and the stability of neutron stars. Specifically, we explored how the properties of these DANS are modified with respect to those of ordinary neutron stars in terms of the DM particle mass $m_{\chi}$ and the fraction of DM present in the star. To model ordinary matter we used a state-of-the-art equation of state of $\beta$-stable nuclear matter obtained using the Brueckner-Hartree-Fock quantum many-body approach starting from two-body and three-body nuclear interactions derived from chiral effective field theory. This EOS, as discussed in Sect. III.1, reproduces various empirical nuclear matter properties at the saturation density $n_{0}$ and at supranuclear densities ($n>\,n_{0}$). Additionally, the calculated $E_{sym}(n_{0})$ and $L$ fulfill the bounds imposed by the unitary Fermi gas limit Tews et al. (2017). The dark matter component of the star has been modeled as a non-self-annihilating system of spin $1/2$ fermions and its EOS as an ideal relativistic Fermi gas. We found that, depending on the dark matter particle mass $m_{\chi}$, one can have somehow opposite effects on the stellar properties. In fact, in the case $m_{\chi}=1\,\mathrm{GeV}$, the stellar gravitational maximum mass $M_{max}$ decreases, whereas in the case $m_{\chi}=0.1\,\mathrm{GeV}$ $M_{max}$ increases with respect to the maximum mass of ordinary neutron stars. We studied $M_{max}$ as a function of $f_{\chi}$ for some values of $m_{\chi}$. Following Ivanytskyi et al. (2020), we defined the critical values $f_{\chi}^{crit}(m_{\chi})$ as the roots of the equation $M(f_{\chi},m_{\chi})=2\ \mathrm{M_{\odot}}$ for a given value of $m_{\chi}$. Comparing an upper limit for the total dark matter fraction accumulated by a $2\ \mathrm{M_{\odot}}$ neutron star in the galactic center with the function $f_{\chi}^{crit}(m_{\chi})$, we found an upper limit for the dark matter particles: $m_{\chi}\leq 56\ \mathrm{GeV}$. We also shown that the presence of dark matter has indirect sizeable effect on the proton fraction in the ordinary matter fluid and, in the case $m_{\chi}=1\,\mathrm{GeV}$, results in a decrease of the threshold gravitational mass $M_{tot}^{durca}$ for having direct URCA processes and fast stellar cooling.

### v2

@@ 7 Summary
In this paper we have studied the consequences of the presence of DM on the structural properties and the stability of neutron stars. Specifically, we explored how the properties of these DANS are modified with respect to those of ordinary neutron stars in terms of the DM particle mass $m_{\chi}$ and the fraction of DM present in the star. To model ordinary matter we used a state-of-the-art equation of state of $\beta$-stable nuclear matter obtained using the Brueckner-Hartree-Fock quantum many-body approach starting from two-body and three-body nuclear interactions derived from chiral effective field theory. This EOS, as discussed in Sect. 3.1, reproduces various empirical nuclear matter properties at the saturation density $n_{0}$ and at supranuclear densities ($n>\,n_{0}$). Additionally, the calculated $E_{sym}(n_{0})$ and $L$ fulfill the bounds imposed by the unitary Fermi gas limit (Tews et al., 2017) and are compatible with the values extracted in Ref. (Reed et al., 2021) from the measured neutron skin thickness of the 208Pb nucleus in the PREX-2 experiment (Adhikari et al., 2021). The dark matter component of the star has been modeled as a non-self-annihilating system of spin $1/2$ fermions and its EOS as an ideal relativistic Fermi gas. We found that, depending on the dark matter particle mass $m_{\chi}$, one can have somehow opposite effects on the stellar properties when a DM fraction $f_{\chi}$ is added to the stellar material. In fact, in the case $m_{\chi}=1\,\mathrm{GeV}$, the stellar gravitational maximum mass $M_{max}$ decreases, whereas in the case $m_{\chi}=0.1\,\mathrm{GeV}$ $M_{max}$ increases with respect to the maximum mass of ordinary neutron stars. We studied $M_{max}$ as a function of $f_{\chi}$ for some fixed values of the DM particle mass $m_{\chi}$. We defined the critical values $f_{\chi}^{crit}(m_{\chi})$ as the roots of the equation $M(f_{\chi},m_{\chi})=2\ \mathrm{M_{\odot}}$ for a given value of $m_{\chi}$. Comparing an upper limit for the total dark matter fraction accumulated by a $2\ \mathrm{M_{\odot}}$ neutron star in the galactic center with the function $f_{\chi}^{crit}(m_{\chi})$, we found an upper limit for the dark matter particles: $m_{\chi}\leq 56\ \mathrm{GeV}$. We also shown that the presence of dark matter has indirect sizable effect on the proton fraction in the ordinary matter fluid and, in the case $m_{\chi}=1\,\mathrm{GeV}$, results in a decrease of the threshold gravitational mass $M_{tot}^{durca}$ for having direct URCA processes and fast stellar cooling.